a man with some passenger in a Boat starts perpendicular to the flow of river 200 m wide and flowing with 2 metre per second boat speed in still water is 4 metre per second when he reaches half the width of the river the passengers asked him they want to reach the just opposite end from where they have started find how many times more total time it would take that if he would have denied the passengers

Grade:11

2 Answers

Vikas TU
4 years ago
Hiii 
Time to cover half of width = 100/4 = 25 sec
shift in direction of flow = 2* 25 = 50 m 
OB = sqrt(100^2 + 50^2)
Time to cover it = 50sqrt 2 / (6/sqrt 5 ) = 41.66
Total time = 25 + 41.66 = 66.66
so, it will take 66.67 /50  more time 
 
 
Khimraj
4 years ago

its a simple sum from dc pandey
initially the vel is perpendicular to the flow so the toal drift is 2*t where t is the time obvousy t is 100/4=25s thereofre =50m
now it has to move in a direction so that the drift becomes 50m in the opp direction
let the anfle be @ with the perpendicular so vel along the horizontal direction is 4sin@ and along the normal 4cos@

time taken to reach the other end is 25/cos@ (simple math)
now drift will be 4sin@-2*25/cos@=50m